plz answer the questions with solutions as asked.

A: If Mariam earned $64,000 in 2015, what should she have been earning in 2018 in order to keep up with inflation if CPI for 2015 is 126.5 and for 2018 is 133.4

B:

Rainbow Retirement Home currently accommodates 250 residents with the staff of 15 personal support workers and 18 nurses. The facility is planning an expansion to allow for accommodation of another 40 residents. Assuming the same proportionate staffing​ levels, how many more nurses will need to be​ hired
Answer #1

A. Earnings in 2015 = $64,000 CPI 2015 = 126.50 CPI 2018 = 133.40

Earning required in 2018 = Earnings 2015/CPI 2015 x CPI 2018 = $64,000/126.5 x 133.40 = $67,491

B.Current Accomodation = 250 residents Current personal staff = 15 Current nurses = 18

Current staffing level personal staff = 250/15 = 16.67

Current staffing level nurses = 250/18 = 13.89

Proposed extra accomodation = 40

Thus extra personal staff = 40 / 16.67 = 2.40

Thus extra nurses = 40 / 13.89 = 2.88 (3 rounded)